OpenStudy (anonymous):

is the sum and the product of a fraction the same thing: example: find the product of 1 1/3 and 5/6? find the sum of 1 1/3 and 5/6? thanks

OpenStudy (anonymous):

sum means "add" and product means "multiply" so no, they are not the same

OpenStudy (anonymous):

sum refers to the addition whereas product means multiplication so the product of : 4/3 * 5/6 = 20/18 and the sum of : 4/3 + 5/6 = 39/18

OpenStudy (anonymous):

note: i converted 1 1/3 to an improper fraction so that its easier to work with

OpenStudy (anonymous):

\[1\tfrac{1}{3}+\frac{5}{6}=\frac{4}{3}+\frac{5}{6}=\frac{8}{6}+\frac{5}{6}=\frac{13}{6}=2\tfrac{1}{6}\]

OpenStudy (anonymous):

whereas \[1\tfrac{1}{3}\times \frac{5}{6}=\frac{4}{3}\times \frac{5}{6}=\frac{20}{18}=1\tfrac{1}{9}\]

OpenStudy (anonymous):

also, when you simplify 39/18 it becomes 13/6